9.A 2-year Treasury security currently earns 4.63percent. Over the next 2 years, the real interest rate is expected to be 2.15percentper year and the inflation premium is expected to be 1.75percentper year. If you make the usual assumption that there is no liquidity risk on a Treasury security, what is the maturity risk premium on the 2-year Treasury security?

A. morethan 2.00 percent

B. more than 1.65 percentbut lessthan 2.00 percent

C. more than 1.20 percentbut lessthan 1.65 percent

D. more than 0.85 percentbut lessthan 1.20 percent

E. lessthan 0.85 percent

Exhibit 1Real interest rate = 1.35percentDefault risk premium = 2.75percentLiquidity risk premium = 0.85percentMaturity risk premium = 1.25percent10.

considering the purchase of 30-year corporate bonds. The bonds have no special covenants. The Wall Street Journal reports that 1-year T-bills are currently earning 2.95 percent. Jane twill use the information listed in Exhibit 1about economic activity and the corporate bonds. What is the value that she calculates for the inflation premium?

A. less than 1.30 percent

B. lessthan 1.55 percentbut more than 1.30 percent

C. lessthan 1.70percentbut more than 1.55 percent

D. lessthan 1.95 percentbut more than 1.70 percent

E. morethan 1.95 percent
Answer #1

9 maturity risk premium = 4.63 - 2.15 - 1.75 = 0.73

E. lessthan 0.85 percent

10 . inflation premium = 2.95 - 1.35 = 1.60

C. lessthan 1.70percentbut more than 1.55 percent